A 90 W electric blanket runs at 24 V .

Part A

What is the resistance of the wire in the blanket?

Express your answer to two significant figures and include the appropriate units.

Part B

How much current does the wire carry?

Express your answer to two significant figures and include the appropriate units.
Answer #1

The resistance is:

\(R=\frac{V^{2}}{P}=\frac{24^{2}}{90}=6.4 \Omega\)

The current is:

\(i=\frac{V}{R}=\frac{24}{6.4}=3.8 \mathrm{~A}\)